Scenario

In this scenario we are going to assume that a factory produces links that are used to make a paperchain. A link is a strip of paper approximately 5 ½” X 1 ¼” cut from an 8 ½” X 11” sheet of paper. The strip is then wrapped into a circle and taped to form a ring. The next strip is cut, placed through the previously made ring and taped to interconnect the 2 paper rings to form the beginnings of a paper chain.

The links factory produced the following results:

When 1 worker was employed, the factory produced 2 links

When 2 workers were employed, the factory produced 6 links

When 3 workers were employed, the factory produced 8 links

When 4 workers were employed, the factory produced 9 links

When 5 workers were employed, the factory produced 9 links

  1. Complete the following table based on the production in the plant.

Workers (Inputs)

Total Product

Marginal Product

0

0

--------

1

2

3

4

5

  1. Did diminishing returns set in in the link factory? At what point?

  1. Explain why diminishing marginal returns sets in at some point in the short run

Consumer and Producer Surplus

Bill buys a computer for $425 for which he was willing to pay $550. The minimum acceptable price to the seller, Beth, was $375.

Bill received _____________________ surplus of _____________.

Beth received _____________________ surplus of ________________.

Diminishing Utility

The first slice of pizza Jake consumers yields him 25 units of utility. The second slice yields him an additional 20 units of utility. His total utility from three slices of pizza is 58 units of utility.

The marginal utility of the third slice of pizza is how much?
Answer #1

Table 1

Workers TP MP
0 0
1 2 2
2 6 4
3 8 2
4 9 1
5 9 0

The marginal product of labor is the change in the total product due to one unit change in labor. Here as labor increases by each unit the amount of chain produce rises. The marginal product is calculated by subtracting the total product of nth labor for the total product of n-1 th labor (previous one).

From the table, it is evident that the diminishing marginal productivity sets in from the 3rd labor.

In the short run, the firms have fixed capital. As the firm keeps on increasing labor the ratio of capital to labor rises. Too much labor depreciates the capital much faster and production increases at a decreasing rate. Hence, in the short run, there is diminishing marginal productivity of labor.

==============================================================================

The consumer surplus is the difference between the price consumer wants to pay and the equilibrium or market price. the producer surplus is the difference between the price the producer willing t receive and the market price. therefore, the correct options are:

  • Consumer; (550-425)=125
  • Producer; (425-375)=50

=============================================================================

The utility from 1st unit is 25, the additional utility from 2nd unit is 20, therefore, the utility from 2nd unit is (25+20)=45. The marginal utility is the additional utility from the last unit consumed. then the marginal utility from 3rd unit is (58-45)=13.

The correct answer is: 13